# Distance measurements from the internal dynamics of globular clusters: Application to the Sombrero galaxy (M 104)

### v1

@@ Appendix A Table of dispersion measurements
We report the velocity dispersions, magnitudes, and where available sizes and mass-to-light ratios in Table 3. The identifiers refer to the catalogue presented in Fahrion et al. (2025).
| ID | RA | DEC | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | |
| 12 (SUCD1) | 190.01304 | 11.66786 | 0.34 | … | 17.461 | … | 3.380.33 | |
| 13 | 190.03900 | 11.64231 | 0.12 | 18.91 | 18.87 | … | 3.930.21 | |
| 14 | 190.04279 | 11.57125 | 0.07 | 20.11 | 20.07 | 20.14 | 1.160.47 | |
| 15 | 190.04479 | 11.78836 | … | … | … | 21.29 | … | |
| 16 | 190.04667 | 11.50436 | … | … | … | 21.21 | … | |
| 17 | 190.04687 | 11.84981 | … | … | … | 20.50 | … | |
| 18 | 190.05292 | 11.40906 | … | … | … | 21.53 | … | |
| 19 | 190.03271 | 11.38036 | … | … | … | 21.04 | … | |
| 20 | 190.05500 | 11.28775 | … | … | … | … | … | |
| 21 | 190.07171 | 11.95222 | … | … | … | … | … | |
| 23 | 190.09050 | 11.81750 | … | … | … | … | … | |
| 24 | 190.09571 | 11.79056 | … | … | … | 20.93 | … | |
| 25 | 190.13242 | 11.93833 | … | … | … | … | … | |
| 26 | 190.06908 | 11.55500 | … | … | … | 21.18 | … | |
| 27 | 190.03042 | 11.43528 | … | … | … | 20.88 | … | |
| 28 | 190.02520 | 11.59367 | 0.06 | 20.44 | 20.53 | 20.55 | 3.140.71 | |
| 29 | 189.93025 | 11.81147 | … | … | … | 20.85 | … | |
| 30 | 189.95262 | 11.59462 | 0.10 | 19.09 | 19.04 | … | 2.710.17 | |
| 33 | 189.97587 | 11.57515 | 0.11 | 19.93 | 19.80 | 19.83 | 3.600.57 | |
| 34 | 189.97846 | 11.68622 | … | … | … | … | … | |
| 35 | 189.98367 | 11.71906 | … | … | … | 20.98 | … | |
| 36 | 190.02600 | 11.85794 | … | … | … | 20.57 | … | |
| 37 | 189.98450 | 11.53192 | … | … | … | … | … | |
| 39 | 190.02096 | 11.48725 | … | … | … | 21.00 | … | |
| 40 | 190.02262 | 11.93269 | … | … | … | … | … | |
| 41 | 190.02500 | 11.49014 | … | … | … | 21.35 | … | |
| 42 | 189.98825 | 11.92911 | … | … | … | … | … | |
| 75 | 189.98917 | 11.64292 | … | … | … | … | … | |
| 91 | 189.99075 | 11.60258 | 0.07 | 19.39 | 19.36 | … | 1.770.21 | |
| 134 | 190.16358 | 11.75308 | … | … | … | 20.70 | … | |
| 135 | 190.16233 | 11.91731 | … | … | … | 19.68 | … | |
| 136 | 190.15450 | 11.76708 | … | … | … | 19.29 | … | |
| 137 | 190.15421 | 11.49092 | … | … | … | 19.44 | … | |
| 139 | 190.15133 | 11.68289 | … | … | … | 20.37 | … | |
| 143 | 190.14529 | 11.47961 | … | … | … | 20.02 | … | |
| 145 | 190.14083 | 11.75581 | … | … | … | 20.45 | … | |
| 154 | 190.11308 | 11.47386 | … | … | … | 20.68 | … | |
| 160 | 190.10167 | 11.89797 | … | … | … | 21.06 | … | |
| 168 | 190.07933 | 11.39892 | … | … | … | 20.25 | … | |
| 172 | 190.07371 | 11.67914 | … | … | … | 20.23 | … | |
| 181 | 190.06475 | 11.60589 | 0.08 | 20.21 | 20.18 | 20.19 | 0.610.33 | |
| 182 | 190.06408 | 11.62633 | 0.11 | 20.76 | 20.79 | 20.62 | 2.950.80 | |
| 184 | 190.06367 | 11.67028 | 0.08 | 19.37 | 19.38 | 19.33 | 3.580.23 | |
| 189 | 190.06083 | 11.53214 | … | … | … | 19.48 | … | |
| 192 | 190.05729 | 11.61739 | 0.04 | 19.80 | 19.89 | 19.91 | 2.080.34 | |
| 194 | 190.05654 | 11.65957 | 0.07 | 19.77 | 19.77 | 19.79 | 1.760.32 | |
| 198 | 190.05179 | 11.42597 | … | … | … | 19.55 | … | |
| 210 | 190.04362 | 11.75575 | … | … | … | 20.31 | … | |
| 211 | 190.04333 | 11.61078 | 0.08 | 19.02 | 18.99 | 18.98 | 1.790.14 | |
| 215 | 190.04037 | 11.60633 | 0.09 | 19.57 | 19.43 | 19.45 | 1.860.20 | |
| 218 | 190.03883 | 11.58786 | 0.06 | 19.75 | 19.83 | 19.85 | 3.510.43 | |
| 221 | 190.03587 | 11.65861 | 0.09 | 19.94 | 19.92 | 19.85 | 2.660.32 |
| ID | RA | DEC | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | |
| 222 | 190.03537 | 11.72353 | … | … | … | 20.38 | … | |
| 226 | 190.03179 | 11.76808 | … | … | … | 20.02 | … | |
| 231 | 190.02949 | 11.60131 | 0.10 | 19.79 | 19.79 | 19.67 | 2.260.22 | |
| 239 | 190.02342 | 11.54158 | … | … | … | 20.38 | … | |
| 240 | 190.02275 | 11.43819 | … | … | … | 20.06 | … | |
| 248 | 190.01504 | 11.48775 | … | … | … | 21.39 | … | |
| 249 | 190.01500 | 11.54497 | … | … | … | 20.48 | … | |
| 255 | 190.00900 | 11.50750 | … | … | … | 19.87 | … | |
| 264 | 190.00225 | 11.74897 | … | … | … | 19.77 | … | |
| 276 | 189.99562 | 11.58962 | 0.08 | 19.51 | 19.49 | 19.53 | 2.010.20 | |
| 279 | 189.99167 | 11.57417 | 0.06 | 20.20 | 20.27 | 20.28 | 2.240.29 | |
| 281 | 189.98996 | 11.52369 | … | … | … | 19.28 | … | |
| 283 | 189.98700 | 11.33994 | … | … | … | 20.25 | … | |
| 288 | 189.98400 | 11.76928 | … | … | … | 20.31 | … | |
| 292 | 189.98221 | 11.65899 | 0.08 | 20.32 | 20.30 | 20.30 | 2.330.29 | |
| 293 | 189.98083 | 11.55678 | … | … | … | 19.81 | … | |
| 302 | 189.97129 | 11.85125 | … | … | … | 20.66 | … | |
| 308 | 189.96717 | 11.63383 | 0.06 | 20.23 | 20.27 | 20.29 | 3.110.52 | |
| 320 | 189.96316 | 11.64266 | 0.11 | 19.92 | 19.96 | 19.70 | 3.230.32 | |
| 339 | 189.94404 | 11.61414 | 0.10 | 19.84 | 19.82 | 19.68 | 2.810.32 | |
| 340 | 189.94371 | 11.58261 | … | … | … | 19.88 | … | |
| 346 | 189.93874 | 11.60014 | 0.13 | 19.33 | 19.28 | 19.13 | 3.670.26 | |
| 355 | 189.92792 | 11.71342 | … | … | … | 20.38 | … | |
| 376 | 189.90033 | 11.56905 | … | … | … | 20.65 | … | |
| 378 | 189.89896 | 11.79611 | … | … | … | 20.61 | … | |
| 381 | 189.89629 | 11.86453 | … | … | … | 19.98 | … | |
| 382 | 189.89621 | 11.50647 | … | … | … | 21.66 | … | |
| 387 | 189.88625 | 11.50875 | … | … | … | 20.18 | … | |
| 395 | 189.87221 | 11.49267 | … | … | … | 19.65 | … | |
| 412 | 190.04184 | 11.63909 | 0.08 | 18.79 | 18.74 | … | 3.260.26 | |
| 413 | 190.00021 | 11.60261 | 0.09 | 18.94 | 18.83 | … | 2.200.30 | |
| 414 | 190.06078 | 11.64111 | 0.11 | 18.98 | 18.83 | … | 2.290.20 | |
| 423 | 189.99700 | 11.64114 | 0.04 | 20.20 | 20.29 | … | 1.970.40 | |
| 424 | 189.96816 | 11.64635 | 0.09 | 20.20 | 20.13 | … | 2.290.61 | |
| 429 | 189.97517 | 11.60142 | 0.07 | 20.44 | 20.48 | … | 2.820.73 |

### v2

@@ Appendix B Table of dispersion measurements
We report the velocity dispersions, magnitudes, and where available sizes, mass-to-light ratios, and metallicities in Table 3. The identifiers refer to the catalogue presented in Fahrion et al. (2025).
| ID1 | RA | DEC | [M/H] | Metal ref2. | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | (dex) | ||
| 12 | 190.01304 | 11.66786 | 0.34 | … | 17.46 | … | 3.380.33 | 0.480.06 | FLAMES | |
| 13 | 190.03900 | 11.64231 | 0.12 | 18.91 | 18.87 | … | 3.930.21 | 0.830.05 | FLAMES | |
| 14 | 190.04279 | 11.57125 | 0.07 | 20.11 | 20.07 | 20.14 | 1.160.47 | 1.760.11 | FLAMES | |
| 15 | 190.04479 | 11.78836 | … | … | … | 21.29 | … | 0.770.17 | FLAMES | |
| 16 | 190.04667 | 11.50436 | … | … | … | 21.21 | … | 1.140.24 | FLAMES | |
| 17 | 190.04687 | 11.84981 | … | … | … | 20.50 | … | 1.490.06 | FLAMES | |
| 18 | 190.05292 | 11.40906 | … | … | … | 21.53 | … | 0.430.23 | FLAMES | |
| 19 | 190.03271 | 11.38036 | … | … | … | 21.04 | … | 0.290.25 | FLAMES | |
| 20 | 190.05500 | 11.28775 | … | … | … | … | … | 0.510.16 | FLAMES | |
| 21 | 190.07171 | 11.95222 | … | … | … | … | … | 1.440.07 | FLAMES | |
| 23 | 190.09050 | 11.81750 | … | … | … | … | … | 0.790.17 | FLAMES | |
| 24 | 190.09571 | 11.79056 | … | … | … | 20.93 | … | 1.790.17 | FLAMES | |
| 25 | 190.13242 | 11.93833 | … | … | … | … | … | 1.750.08 | FLAMES | |
| 26 | 190.06908 | 11.55500 | … | … | … | 21.18 | … | 0.510.32 | FLAMES | |
| 27 | 190.03042 | 11.43528 | … | … | … | 20.88 | … | 2.150.15 | FLAMES | |
| 28 | 190.02520 | 11.59367 | 0.06 | 20.44 | 20.53 | 20.55 | 3.140.71 | 0.940.08 | OSIRIS | |
| 29 | 189.93025 | 11.81147 | … | … | … | 20.85 | … | 1.310.08 | FLAMES | |
| 30 | 189.95262 | 11.59462 | 0.10 | 19.09 | 19.04 | … | 2.710.17 | 0.820.09 | OSIRIS | |
| 33 | 189.97587 | 11.57515 | 0.11 | 19.93 | 19.80 | 19.83 | 3.600.57 | 1.080.12 | OSIRIS | |
| 34 | 189.97846 | 11.68622 | … | … | … | … | … | 1.140.41 | FLAMES | |
| 35 | 189.98367 | 11.71906 | … | … | … | 20.98 | … | 1.480.25 | FLAMES | |
| 36 | 190.02600 | 11.85794 | … | … | … | 20.57 | … | 1.690.14 | FLAMES | |
| 37 | 189.98450 | 11.53192 | … | … | … | … | … | 0.000.08 | FLAMES | |
| 39 | 190.02096 | 11.48725 | … | … | … | 21.00 | … | 0.400.07 | FLAMES | |
| 40 | 190.02262 | 11.93269 | … | … | … | … | … | 1.790.03 | FLAMES | |
| 41 | 190.02500 | 11.49014 | … | … | … | 21.35 | … | 1.490.17 | FLAMES | |
| 42 | 189.98825 | 11.92911 | … | … | … | … | … | 1.330.09 | FLAMES | |
| 75 | 189.98917 | 11.64292 | … | … | … | … | … | 1.850.20 | MUSE | |
| 91 | 189.99075 | 11.60258 | 0.07 | 19.39 | 19.36 | … | 1.770.21 | 1.360.13 | MUSE | |
| 134 | 190.16358 | 11.75308 | … | … | … | 20.70 | … | 1.400.13 | FLAMES | |
| 135 | 190.16233 | 11.91731 | … | … | … | 19.68 | … | 1.070.13 | FLAMES | |
| 136 | 190.15450 | 11.76708 | … | … | … | 19.29 | … | 0.980.09 | FLAMES | |
| 137 | 190.15421 | 11.49092 | … | … | … | 19.44 | … | 1.060.15 | FLAMES | |
| 139 | 190.15133 | 11.68289 | … | … | … | 20.37 | … | 1.260.05 | FLAMES | |
| 143 | 190.14529 | 11.47961 | … | … | … | 20.02 | … | 1.060.14 | FLAMES | |
| 145 | 190.14083 | 11.75581 | … | … | … | 20.45 | … | 0.930.12 | FLAMES | |
| 154 | 190.11308 | 11.47386 | … | … | … | 20.68 | … | 0.710.16 | FLAMES | |
| 160 | 190.10167 | 11.89797 | … | … | … | 21.06 | … | 1.490.03 | FLAMES | |
| 168 | 190.07933 | 11.39892 | … | … | … | 20.25 | … | 0.920.12 | FLAMES | |
| 172 | 190.07371 | 11.67914 | … | … | … | 20.23 | … | 1.260.00 | FLAMES | |
| 181 | 190.06475 | 11.60589 | 0.08 | 20.21 | 20.18 | 20.19 | 0.610.33 | 1.120.13 | FLAMES | |
| 182 | 190.06408 | 11.62633 | 0.11 | 20.76 | 20.79 | 20.62 | 2.950.80 | 1.260.00 | FLAMES | |
| 184 | 190.06367 | 11.67028 | 0.08 | 19.37 | 19.38 | 19.33 | 3.580.23 | 0.910.05 | FLAMES | |
| 189 | 190.06083 | 11.53214 | … | … | … | 19.48 | … | 0.470.15 | FLAMES | |
| 192 | 190.05729 | 11.61739 | 0.04 | 19.80 | 19.89 | 19.91 | 2.080.34 | 0.690.14 | FLAMES | |
| 194 | 190.05654 | 11.65957 | 0.07 | 19.77 | 19.77 | 19.79 | 1.760.32 | 1.120.23 | OSIRIS | |
| 198 | 190.05179 | 11.42597 | … | … | … | 19.55 | … | 0.540.12 | FLAMES | |
| 210 | 190.04362 | 11.75575 | … | … | … | 20.31 | … | 0.430.17 | FLAMES | |
| 211 | 190.04333 | 11.61078 | 0.08 | 19.02 | 18.99 | 18.98 | 1.790.14 | 0.720.07 | FLAMES | |
| 215 | 190.04037 | 11.60633 | 0.09 | 19.57 | 19.43 | 19.45 | 1.860.20 | 1.280.09 | FLAMES | |
| 218 | 190.03883 | 11.58786 | 0.06 | 19.75 | 19.83 | 19.85 | 3.510.43 | 0.570.10 | FLAMES | |
| 221 | 190.03587 | 11.65861 | 0.09 | 19.94 | 19.92 | 19.85 | 2.660.32 | 0.760.08 | FLAMES |
| ID | RA | DEC | [M/H] | Metal ref. | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | (dex) | ||
| 222 | 190.03537 | 11.72353 | … | … | … | 20.38 | … | 1.360.12 | FLAMES | |
| 226 | 190.03179 | 11.76808 | … | … | … | 20.02 | … | 0.900.11 | FLAMES | |
| 231 | 190.02949 | 11.60131 | 0.10 | 19.79 | 19.79 | 19.67 | 2.260.22 | 0.160.06 | OSIRIS | |
| 239 | 190.02342 | 11.54158 | … | … | … | 20.38 | … | 1.130.18 | FLAMES | |
| 240 | 190.02275 | 11.43819 | … | … | … | 20.06 | … | 0.500.21 | FLAMES | |
| 248 | 190.01504 | 11.48775 | … | … | … | 21.39 | … | 0.240.29 | FLAMES | |
| 249 | 190.01500 | 11.54497 | … | … | … | 20.48 | … | 0.100.14 | FLAMES | |
| 255 | 190.00900 | 11.50750 | … | … | … | 19.87 | … | 0.340.11 | FLAMES | |
| 264 | 190.00225 | 11.74897 | … | … | … | 19.77 | … | 0.210.12 | FLAMES | |
| 276 | 189.99562 | 11.58962 | 0.08 | 19.51 | 19.49 | 19.53 | 2.010.20 | 0.930.09 | OSIRIS | |
| 279 | 189.99167 | 11.57417 | 0.06 | 20.20 | 20.27 | 20.28 | 2.240.29 | 0.270.11 | FLAMES | |
| 281 | 189.98996 | 11.52369 | … | … | … | 19.28 | … | 0.400.01 | FLAMES | |
| 283 | 189.98700 | 11.33994 | … | … | … | 20.25 | … | 1.410.17 | FLAMES | |
| 288 | 189.98400 | 11.76928 | … | … | … | 20.31 | … | 1.490.04 | FLAMES | |
| 292 | 189.98221 | 11.65899 | 0.08 | 20.32 | 20.30 | 20.30 | 2.330.29 | 0.270.17 | OSIRIS | |
| 293 | 189.98083 | 11.55678 | … | … | … | 19.81 | … | 1.140.10 | FLAMES | |
| 302 | 189.97129 | 11.85125 | … | … | … | 20.66 | … | 1.110.15 | FLAMES | |
| 308 | 189.96717 | 11.63383 | 0.06 | 20.23 | 20.27 | 20.29 | 3.110.52 | 1.050.23 | FLAMES | |
| 320 | 189.96316 | 11.64266 | 0.11 | 19.92 | 19.96 | 19.70 | 3.230.32 | 0.240.05 | OSIRIS | |
| 339 | 189.94404 | 11.61414 | 0.10 | 19.84 | 19.82 | 19.68 | 2.810.32 | 0.050.10 | FLAMES | |
| 340 | 189.94371 | 11.58261 | … | … | … | 19.88 | … | 1.000.11 | FLAMES | |
| 346 | 189.93874 | 11.60014 | 0.13 | 19.33 | 19.28 | 19.13 | 3.670.26 | 0.160.04 | OSIRIS | |
| 355 | 189.92792 | 11.71342 | … | … | … | 20.38 | … | 1.490.06 | FLAMES | |
| 376 | 189.90033 | 11.56905 | … | … | … | 20.65 | … | 1.870.15 | OSIRIS | |
| 378 | 189.89896 | 11.79611 | … | … | … | 20.61 | … | 1.260.12 | FLAMES | |
| 381 | 189.89629 | 11.86453 | … | … | … | 19.98 | … | 1.100.11 | FLAMES | |
| 382 | 189.89621 | 11.50647 | … | … | … | 21.66 | … | 0.260.12 | FLAMES | |
| 387 | 189.88625 | 11.50875 | … | … | … | 20.18 | … | 0.690.13 | FLAMES | |
| 395 | 189.87221 | 11.49267 | … | … | … | 19.65 | … | 1.200.11 | FLAMES | |
| 412 | 190.04184 | 11.63909 | 0.08 | 18.79 | 18.74 | … | 3.260.26 | 0.820.10 | OSIRIS | |
| 413 | 190.00021 | 11.60261 | 0.09 | 18.94 | 18.83 | … | 2.200.30 | 0.880.10 | MUSE | |
| 414 | 190.06078 | 11.64111 | 0.11 | 18.98 | 18.83 | … | 2.290.20 | 0.570.10 | OSIRIS | |
| 423 | 189.99700 | 11.64114 | 0.04 | 20.20 | 20.29 | … | 1.970.40 | 0.220.08 | MUSE | |
| 424 | 189.96816 | 11.64635 | 0.09 | 20.20 | 20.13 | … | 2.290.61 | 1.010.13 | OSIRIS | |
| 429 | 189.97517 | 11.60142 | 0.07 | 20.44 | 20.48 | … | 2.820.73 | 0.360.23 | MUSE |